OpenStudy (anonymous):

A 14 foot ladder is leaning against a building. The ladder makes a 45 degree angle with the building. How far up the building does the ladder reach?

OpenStudy (faiqraees):



rishavraj (rishavraj):

\[clue~~\sin \theta = \frac{ opposite }{ hypotenuse }\]

OpenStudy (anonymous):

What does the bottom line equal?

OpenStudy (faiqraees):

You can ignore that

OpenStudy (anonymous):

Does anyone know the answer?

OpenStudy (faiqraees):

yes

OpenStudy (anonymous):

What is it?

OpenStudy (snowsurf):

 Find h. @rishavraj already gave you a clue.